OpenStudy (anonymous):

can someone help me understand the quadratic formula and the discriminant.. with an example please :)

OpenStudy (amistre64):

you ever complete a square?

OpenStudy (amistre64):

the quadratic formula is the general result of completing the square .... as such, instead of working the process in a longhand version, we can simply go straight to the ending form of it.

OpenStudy (anonymous):

The discriminant is meant to tell you if the solutions are distinct, repeated, or imaginary.

OpenStudy (amistre64):

\[ax^2+bx+c=0\] \[x^2+\frac bax+\frac ca=0\] \[x^2+\frac bax=-\frac ca\] \[x^2+\frac bax+(\frac{b}{2a})^2=(\frac{b}{2a})^2-\frac ca\] \[(x+\frac{b}{2a})^2=(\frac{b}{2a})^2-\frac ca\] \[x+\frac{b}{2a}=\pm\sqrt{(\frac{b}{2a})^2-\frac ca}\] \[x=-\frac{b}{2a}\pm\sqrt{(\frac{b}{2a})^2-\frac ca}\]

OpenStudy (amistre64):

with a little more simplification you would see that the quadrtic formula is the end result of the gneneral solution to the quadratic equation.

OpenStudy (anonymous):

@iambatman

OpenStudy (amistre64):

now if we simply ues the formula, given ax^2 + bx+c=0 \[x=-\frac{b}{2a}\pm\frac{\sqrt{b^2-4ac}}{2a}\] just fill in the required parts for abc

OpenStudy (amistre64):

make what you have into the form: ax^2 + bx + c = 0 so that you can identify the required parts

OpenStudy (amistre64):

the quadratic formula is so useful that they have even named parts of it -b/2a is called the determinant, and b^2-4ac is called the discriminant. -b/2a tells us the axis of symetry, and the b^2-4ac lets us know wht types of roots we have (real or complex)